what are the solutions to x^2+4x-12=0
Answer by haileytucki(390)

x^(2)+4x-12=0
In this problem 6*-2=-12 and 6-2=4, so insert 6 as the right hand term of one factor and -2 as the right-hand term of the other factor.
(x+6)(x-2)=0
Set each of the factors of the left-hand side of the equation equal to 0.
x+6=0_x-2=0
Since 6 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 6 from both sides.
x=-6_x-2=0
Set each of the factors of the left-hand side of the equation equal to 0.
x=-6_x-2=0
Since -2 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 2 to both sides.
x=-6_x=2
The complete solution is the set of the individual solutions.
x=-6,2